------------------------------------------------------------
Date: Tue Oct 2 00:01:43 2001, Shift: 16:00-24:00
Supervisor: R. Debbe, with: none
Log book#9 pages 71 - 73
Shift Summary:
------------------------------------------------------------
Machine status: Good store started at 18:02 with 433 ZDC coinc
per second.
Achievements: Took data during most of the shift:
5080 MRS at 60 FS at 3 390K trig 1
5081 D3 and D4 off for alignement 115K
5082 all magnets back on 179K
two sets of pedestal and DVM runs.
Detector Problems: None, I started printing some quality plots
for TPC pedestal, same plots obtained with XTPI should should
be printed before installing new pedestal values. I also estimated
that 3K trigger 7 with DVM data give us good information
on T1,T2 and TPM2 every half an hour.
------------------------------------------------------------
This archive was generated by hypermail 2b30 : Tue Oct 02 2001 - 00:02:30 EDT